Ming Lee spent $10.20 for a pizza and a soft drink. If the pizza cost 11 times as much as the beverage, how much was each?
Answer by checkley77(12569)

x+11x=10.20
12x=10.20
x=10.20/12
x=$.85 is the price of the soda.
11*.85=9.35 cost of the pizza.
proof:
.85+9.35=1020
1020=10.20